Intermediate Algebra — MAT1001
In a certain triangle the measure of one angle is double the measure of a second angle but is 5 degrees less than the measure of the third angle. If the sum of the measures of the three interior angles of a triangle is always 180 degrees, form an algebraic equation to express the problem, and identify the variables, coefficients, and constants of the algebraic expression.

Answer by stanbon(75887)

In a certain triangle the measure of one angle is double the measure of a second angle but is 5 degrees less than the measure of the third angle.
1st angle = x
2nd angle = 2x
3rd angle 2x+5
---------------------
If the sum of the measures of the three interior angles of a triangle is always 180 degrees, form an algebraic equation to express the problem, and identify the variables, coefficients, and constants of the algebraic expression.
EQUATION:
x + 2x + 2x+5 = 180
5x + 5 = 180
5x = 175
x = 35 degrees
2x = 70 degrees
2x+5 = 75 degrees
=====================
Cheers,
Stan H.
------------------------------------